OpenStudy (anonymous):

What is the slope of a line parallel to the line with equation 5x + 3y = 7?

OpenStudy (anonymous):

@Luigi0210

OpenStudy (luigi0210):

Oh, now you ask for my help >_> But put it into slope-intercept form, and look for m \[\LARGE y=mx+b\]

OpenStudy (anonymous):

But the sentence doesn't even have a y

OpenStudy (luigi0210):

\[\LARGE 5x+3\color{red}{y}=7\]?

OpenStudy (anonymous):

..... o

OpenStudy (anonymous):

What about m and b?

OpenStudy (luigi0210):

Those are just variables representing constants..

OpenStudy (a1234):

5x + 3y = 7 Parallel lines have the same slope, but different y-intercepts. Solve for y: 3y = -5x + 7 Divide by 3: y = -5/3x + 7/3 The slope of the parallel line will be -5/3.

OpenStudy (luigi0210):

^^Pretty much that, even if they did break the CoC >.>

OpenStudy (a1234):

I explained how to do it. It wasn't just a straight answer.

OpenStudy (anonymous):

@Luigi0210